OpenStudy (anonymous):

PLEASE help: Simplify these expressions. 1. √12x^5 · √6x My Answer: I got 6x^3√2. 2. √15/2y^2 3. (6 + √6)(6 - 2√6)

OpenStudy (radar):

\[\sqrt{72x ^{6}}=6x ^{3}\sqrt{2}\]

OpenStudy (radar):

\[36+6\sqrt{6}-12\sqrt{6}-12=24-6\sqrt{6}=6(4-\sqrt{6})\]

OpenStudy (radar):

for 2. is the \[2y ^{2} within the radical?\]

OpenStudy (anonymous):

Yes . Btw, thank you!

OpenStudy (radar):

\[\sqrt{15\over2y ^{2}}=\sqrt{15}/\sqrt{2}y\]